OpenStudy (anonymous):

2(1-2x)/3 - 2x = -2/5 + 4(2-3x)/3 solve the following linear equation - where do i start? thankyou!

OpenStudy (campbell_st):

so just to check the question is \[\frac{2(1 - 2x)}{3} - 2x = -\frac{2}{5} +\frac{ 4(2 - 3x)}{3}\]

OpenStudy (anonymous):

yes!

OpenStudy (cwrw238):

the first thing to do is to get rid of the fractions by multiplying each term by the LCM ( the Lowest Common Multiple) of 3 and 5

OpenStudy (cwrw238):

LCM of 3 and 5 is 15 - can you proceed form here?

OpenStudy (anonymous):